# FaceQSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v2

@@ 4.4.7 Compare Tracker
Table 8: Compare FaceQSORT with state-of-the-art multi-face tracker, where MC and IoU denote FaceQSORT wihtout matching cascade (MC) and/or IoU fallback matching.
@@
FaceQSORT 0.6527 0.7486 0.7338 0.0104 0.8404 0.8672 0.8783 0.0081 0.0425 0.1821 0.0918 0.0343
OMTMCM[42] 0.2039 0.3765 0.3352 0.1671 0.2430 0.4483 0.4140 0.1335 0.1077 0.2842 0.2383 0.0944
MCFDR[36] 0.4770 0.5886 0.0740 0.0097 0.5514 0.5432 0.0144 0.0014 0.0274 0.1142 0.0063 0.0053
UMA[18] 0.1854 0.2531 0.2640 0.0437 0.5483 0.6932 0.6788 0.0112 0.0490 0.1933 0.0988 0.0268
@@
In table 8 FaceQSORT is compared with different state-of-the-art multi-face trackers. A comparison with StrongSORT including MC (i.e., when $\lambda$ is equal 1.0 or 0.0) is already performed in ‘1) Selection of a fixed $\lambda$’. The FaceQSORT results are obtained with SFace features, $\lambda=0.1$ and $\theta=0.2$. Implementations are available for the MCFDR [^7] https://github.com/yjwong1999/OpenVINO-Face-Tracking-using-YOLOv8-and-DeepSORT and UMA [^8] https://github.com/yinjunbo/UMA-MOT. However, no implementation is available for the OMTMCM multi-face tracker. For this reason, OMTMCM is carefully implemented according to the description in [42]. To verify the implementation, the results for the MusicVideo dataset [70] are also presented in table 8. The MusicVideo dataset consists of 8 music videos (i.e., Apink, BrunoMars, Darling, GirlsAloud, HelloBubble, PussycatDolls, Tara and Westlife) from YouTube, is publicly available [^9] https://sites.google.com/site/shunzhang876/eccv16-face-tracking and was used in [42] to evaluate the proposed OMTMCM multi-face tracker. Since music videos are unconstrained videos (with different scenes, moving camera, etc.), they do not correspond to the considered scenario in which people move towards a gate for which FaceQSORT is designed. For this dataset, OMTMCM performed best. However, in the considered sceneraio where people are moving towards a gate (PLUSFiaQ and ChokePoint), FaceQSORT clearly outperformed the other evaluated state-of-the-art trackers.
5 Ablation Study
The main components of FaceQSORT are the combination of biometric and appearance features, the matching cascade and IoU fallback matching. Running FaceQSORT with either biometric (i.e., $\lambda=1.0$) or appearance features (i.e., $\lambda=0.0$) has already been evaluated in ‘1) Selection of a fixed $\lambda$’ of subsection 4.4.
Table 9: Frames per second for FaceQSORT (FS), FaceQSORT without the matching cascade (FS MC) and OMTMCM [42].
PLUSFiaQ ChokePoint
FS FS MC OMTMCM FS FS MC OMTMCM
90.43 87.25 71.14 143.03 134.99 134.01
In the matching cascade (described in section 2), the tracks that were successfully matched in the previous frame are considered first. This can reduce the search space $\Phi$ and can be particularly beneficial if long occlusions are allowed (i.e., a large active track memory $\Phi$). For all experiments, occlusions up to 100 frames are allowed (i.e., $N_{max}=100$). In table 9, execution times in frames per second (fps) are reported for FaceQSORT with and without the matching cascade (MC) and compared with the re-implementation of a state-of-the-art tracker (i.e., OMTMCM [42]). All times were recorded on a standard desktop PC (i.e., Intel Core i5 gen9) and without taking face detection and feature extraction into account. It can be seen that the fps increases slightly when MC is applied. This is also reflected by the observation that on average less than half of all stored active tracks are required to successfully match all detected faces. However, applying the MC carries the risk that not the best matches are found. For example, it could happen that the best match (lowest cost) for a detected face is with a track that could not be matched in the last frames (e.g., a person returning from an occlusion), but the detected face is matched with a track that is tried in the MC before. The achieved tracking performance scores for FaceQSORT (with SFace features, $\lambda=0.1$ and $\theta=0.2$) without the matching cascade (MC) are reported in table 8. Compared to results when using the matching cascade (table 8 first row) a performance decrease can be observed.
@@ 5 Ablation Study
IoU fallback matching is applied to all unmatched detections after the matching cascade, including matching with tentative tracks (only confirmed tracks are considered in the matching cascade). To achieve the reported tracking performance of FaceQSORT (table 8 first row), IoU matching is performed for 3.91%, 4.9% and 8.42% of all detected faces of the PLUSFiaQ, ChokePoint and MusicVideo dataset respectively. When disabling IoU fallback matching (IoU) the performance decreases significantly (see table 8). The performance achieved when FaceQORT is applied without matching cascade and IoU fallback matching is reported in table 8 (last row).
@@ 6 Conclusion
It this work, a new multi-face tracking method, FaceQSORT, is proposed. To mitigate the problem of partially occluded and lateral faces, two different features (i.e., face (biometric) and appearance features) are combined. In the considered scenario, when people move towards a gate/portal (PLUSFiaQ and ChokePoint dataset) FaceQSORT clearly outperformed the evaluated state-of-the-art tracker. To get a deeper insight into the proposed method, a comprehensive experimental evaluation and an ablation study are conducted. It is shown that the selection of the face descriptor and the similarity measure, as well as the resulting distribution of genuine and imposter scores are crucial. In general, selecting a low similarity threshold $\theta$ and a high parameter $\lambda$ lead to a high AssPr.

### v3

@@ 6.2 Tracker Comparison
Table 6: Compare FaceQSORT with state-of-the-art tracker, where MC and IoU denote FaceQSORT without matching cascade (MC) and/or IoU fallback matching. StrongSORT and DeepSORT are applied either with biometric (BIO-SF01) or with appearance features.
@@
FaceQSORT 0.6527 0.7486 0.7338 0.0104 0.8404 0.8671 0.8771 0.0090 0.0425 0.1821 0.0918 0.0343
StrongSORTAPP[21] 0.4825 0.6303 0.6172 0.0177 0.6626 0.7657 0.7729 0.0158 0.0407 0.1720 0.0841 0.0363
StrongSORTBIO[21] 0.4780 0.6256 0.5678 0.0198 0.7300 0.8080 0.7750 0.0099 0.0305 0.1486 0.0637 0.0328
DeepSORTAPP[20] 0.5556 0.6805 0.6769 0.0094 0.7925 0.8418 0.8453 0.0075 0.0429 0.1769 0.0914 0.0287
DeepSORTBIO[20] 0.4700 0.6201 0.5570 0.0191 0.7278 0.8063 0.7740 0.0112 0.0306 0.1488 0.0634 0.0325
DeepOCSORT[64] 0.4816 0.6114 0.5819 0.0172 0.7841 0.8244 0.8225 0.0069 0.0354 0.1585 0.0774 0.0318
OCSORT[65] 0.5030 0.6292 0.6006 0.0143 0.8280 0.8487 0.8502 0.0051 0.0362 0.1598 0.0764 0.0310
BoTSORT[66] 0.4206 0.5257 0.4880 0.0265 0.7092 0.6959 0.6760 0.0139 0.0301 0.1397 0.0614 0.0366
SORT[19] 0.5250 0.6558 0.6006 0.0148 0.8583 0.8804 0.8612 0.0051 0.0336 0.1543 0.0651 0.0312
OMTMCM[36] 0.2039 0.3763 0.3352 0.1671 0.2430 0.4483 0.4140 0.1335 0.1076 0.2841 0.2383 0.0944
UMA[23] 0.1828 0.2475 0.2640 0.0437 0.5481 0.6929 0.6788 0.0112 0.0490 0.1930 0.0988 0.0268
ByteTrack[67] 0.4846 0.6163 0.5831 0.0145 0.8320 0.8509 0.8520 0.0046 0.0349 0.1576 0.0740 0.0287
@@
In table 6 FaceQSORT is compared with different state-of-the-art trackers. The FaceQSORT results are obtained with BIO-SF01 features, $\lambda=0.1$ and $\theta=0.2$. BIO-SF01 features are selected because these features perform best in the considered scenario. StrongSORT and DeepSORT are also evaluated with BIO-SF01 or appearance features (i.e., StrongSORTAPP, StrongSORTBIO, DeepSORTAPP and DeepSORTAPP). If possible, similar parameters (e.g., $\theta$, $N_{max}=100$, etc.) are set identically for all trackers evaluated.
@@
Implementations are available for SORT [^7] https://github.com/abewley/sort, DeepSORT [^8] https://github.com/nwojke/deep_sort, StrongSORT [^9] https://github.com/dyhBUPT/StrongSORT and UMA [^10] https://github.com/yinjunbo/UMA-MOT. A multi-face tracking version of DeepOCSORT [64], OCSORT [65], BoTSORT [66] and ByteTrack [67] with SphereFace features and yolov8 face detection is available in [^11] https://github.com/yjwong1999/OpenVINO-Face-Tracking-using-YOLOv8-and-DeepSORT [68]. No implementation is available for the OMTMCM multi-face tracker. For this reason, OMTMCM is carefully implemented according to the description in [36]. To verify the implementation, the results for the MusicVideo dataset [9] are also presented in table 6. The MusicVideo dataset consists of 8 music videos (i.e., Apink, BrunoMars, Darling, GirlsAloud, HelloBubble, PussycatDolls, Tara and Westlife) from YouTube, is publicly available [^12] https://sites.google.com/site/shunzhang876/eccv16-face-tracking and was used in [36] to evaluate the proposed OMTMCM multi-face tracker. Since music videos are unconstrained videos (with different scenes, moving camera, etc.), they do not correspond to the considered scenario in which people move towards a gate for which FaceQSORT is designed.
For the MusicVideo dataset, OMTMCM performed best. However, in the considered scenario where people are moving towards a gate (PLUSFiaQ), FaceQSORT clearly outperformed the other evaluated state-of-the-art trackers. With the ChokePoint dataset, SORT achieved the highest AssA and HOTA scores. The high performance of SORT with the ChokePoint dataset can be explained by the fact that in the ChokePoint dataset, people walk through the portal with constant (linear) motion. However, FaceQSORT achieved a similarly high AssA and HOTA score and a higher IDF1 score.
6.3 Ablation Study
The main components of FaceQSORT are the combination of biometric and appearance features, the matching cascade (MC) and IoU fallback matching. Running FaceQSORT with either biometric (i.e., $\lambda=1.0$) or appearance features (i.e., $\lambda=0.0$) has already been evaluated in ‘1) Selection of a fixed $\lambda$’ of subsection 6.1. Applying the MC carries the risk that not the best matches are found. For example, it could happen that the best match (lowest cost) for a detected face is with a track that could not be matched in the last frames (e.g., a person returning from an occlusion), but the detected face is matched with a track that is considered in the MC before. The achieved tracking performance scores for FaceQSORT (with BIO-SF01 features, $\lambda=0.1$ and $\theta=0.2$) without the matching cascade (MC) are reported in table 6. Compared to results when using the matching cascade (table 6 first row) a significant performance drop can be observed.
@@ 6.3 Ablation Study
IoU fallback matching is applied to all unmatched detections after the matching cascade, including matching with tentative tracks (only confirmed tracks are considered in the matching cascade). To achieve the reported tracking performance of FaceQSORT (table 6 first row), IoU matching is performed for 3.91%, 4.9% and 8.42% of all detected faces of the PLUSFiaQ, ChokePoint and MusicVideo dataset respectively. When disabling IoU fallback matching (IoU) the performance decreases significantly (see table 6). The performance achieved when FaceQSORT is applied without matching cascade and IoU fallback matching is reported in table 6 (last row).
@@ 7 Conclusion
In this work, a new multi-face tracking method, FaceQSORT, is proposed. To mitigate tracking challenges, two different features (i.e., biometric and appearance features) extracted from the same image (face) patch are combined. It is shown that a combination of the two features is beneficial for multi-face tracking compared to using only one feature type. In the considered scenario, when people move towards a gate (PLUSFiaQ) FaceQSORT clearly outperformed the evaluated state-of-the-art tracker. To get a deeper insight into the proposed method, a comprehensive experimental evaluation and an ablation study are conducted. It is shown that the selection of the face recognition model, as well as the resulting distribution of genuine and imposter scores are crucial. In general, selecting a low similarity threshold $\theta$ and a high parameter $\lambda$ lead to a high AssPr.